OpenStudy (anonymous):

The measure of an angle is one fith the measure of its supplement find the measure of the angle

OpenStudy (mayankdevnani):

@hellokitty2636 lets say one angle be x and other angle be y. So, they are supplementary angles, \[\huge \angle x+\angle y=180 ^ \circ\]

OpenStudy (mayankdevnani):

and it is given that :- The measure of an angle is one fifth the measure of its supplement. so, x's supplement is 180-x. we get, \[\huge x=\frac{1}{5}(180-x)\]

OpenStudy (mayankdevnani):

now solve for x? @hellokitty2636

OpenStudy (aivantettet26):

how do you do that bro? xD

OpenStudy (aivantettet26):

that pics XD

OpenStudy (aivantettet26):

1/5 of 180 is equal to 36 180 - 36 = 144 Not really sure tho

OpenStudy (mayankdevnani):

lol....its 180 - x

OpenStudy (mayankdevnani):

can you solve it? @hellokitty2636

OpenStudy (anonymous):

Thks so much u guys r saving me from my final but I gt more questions

OpenStudy (mayankdevnani):

you are welcome :)

OpenStudy (goldphenoix):

@aivantettet26 , Mayan used the magic equation button on the bottom.

OpenStudy (aivantettet26):

wow. haha now i get it haha

OpenStudy (mayankdevnani):

use `frac` to write in ratio. Like :- `frac{5}{2}`=\(\large \frac{5}{2}\) for this i use :- `\(\large \frac{5}{2}\)`

OpenStudy (mayankdevnani):

@aivantettet26

OpenStudy (goldphenoix):

I told him that already, in private. :)

OpenStudy (mayankdevnani):

lolzz...